OpenStudy (anonymous):

What makes this a perfect square trinomial? x^2+22x+? Is the ? 11

OpenStudy (anonymous):

because 11*2 is 22

OpenStudy (anonymous):

11^2

OpenStudy (anonymous):

Wait the answer is 11^2? But wouldn't that be 121?

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\] and here a = x, 2ab = 22x making b 11, and therefor you want \[x^2+22x+121=(x+11)^2\]\]

OpenStudy (anonymous):

\[?=121\] yes

OpenStudy (anonymous):

11. yes

OpenStudy (anonymous):

????

OpenStudy (anonymous):

this is not an equation. the question was what should the constant be to make the expression a perfect square. the answer is the constant should be 121 so you will have a perfect square, namely \[(x+11)^2=x^2+22x+121\]